OpenStudy (anonymous):

(-7,3) slope= -1/4 Find slope intercept form and standard form

OpenStudy (anonymous):

3= -1/4(-7) + ? what would the y intercept be? -7/3?

OpenStudy (anonymous):

3=-1/4x+b

OpenStudy (jdoe0001):

(y- 3) = -1/4(x-(-7)) solve for "y"

OpenStudy (anonymous):

y-3=-7/4

OpenStudy (anonymous):

sorry: y-3=-1/4 + -7/4

OpenStudy (jdoe0001):

y-3=-1/4 x+ -7/4

OpenStudy (jdoe0001):

hmmm well (y- 3) = -1/4(x-(-7)) => y - 3 = -1/4x +7/4 => y = -1/4x +7/4 +3 y = -1/4x +19/4

OpenStudy (jdoe0001):

ohhh wait shoot

OpenStudy (jdoe0001):

you're correct is -7, - * - + * - = -

OpenStudy (jdoe0001):

so, y = -1/4x +5/4

OpenStudy (anonymous):

so y=-1/4x+5/4 ? because you add 3 to each side?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

ok so what is the y intercept?

OpenStudy (jdoe0001):

y=-1/4x +5/4 ^ ^ slope y-intercept

OpenStudy (anonymous):

thank you so much! i know how to do the rest of the problem now.

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

3=-1/4x+5/4